The demand function for a product is: Qd = 1,000-10P. and its supply function is: Qs = 100 + 2P

The equilibrium price of the good is equal to ____

The equilibrium quantity of the good is equal to ______
Answer #1

Ans:

The equilibrium price of the good is equal to $75

The equilibrium quanity of the good is equal to 250 units.

Explanation:

At equilibrium, Qd = Qs

1000 - 10P = 100 + 2P

12P = 900

P = 900 / 12

= 75

Now put this price either in demand function or in supply function.

Qd =1000 - 10P

= 1000 - ( 10 × 75 )

= 1000 - 750

=250